OpenStudy (anonymous):

Complete the solution of the equation.Find the value of y when x equals 10. 5x-y=54

OpenStudy (anonymous):

y=5x-54 y=5(10)-54 y=-4

OpenStudy (anonymous):

complete the solution of the equation. FInd the value of y when x equals 15. 3x-8y=5

OpenStudy (anonymous):

@damoss

OpenStudy (anonymous):

do the same thing 3(15)-8y=5 y=5